how do you solve for Y in 2x+y=7
Answer by AnlytcPhil(1810)

how do you solve for y in  2x + y = 7

The idea is to get the equation so that it 
starts off like this: "y = "       

                             2x + y = 7

Subtract 2x from both sides:

                             2x + y = 7
                           - 2x         - 2x
                           -----------------
                                  y = 7 - 2x

That's it!

Edwin